OpenStudy (anonymous):

Suppose that f"' is continuous and f'(c) = f"(c) = 0, but f"'(c)>0. Does f have a local maximum or minimum at c? Does f have a point of inflection at c?

OpenStudy (nenadmatematika):

well, take the function y=x^3 for example. y'=3x^2 and y''=6x At the point x=0 y' and y'' have the same value 0